Circular No. 7047

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998S IN NGC 3877 P. Garnavich, P. Challis, and R. Kirshner, Harvard-Smithsonian Cen-

ter for Astrophysics, on behalf of the Supernova Intensive Study collaboration, report that spectra (range 115–570 nm) of SN 1998S were obtained on Oct. 26 with the Space Telescope Imaging Spectrograph (using three gratings with an average resolving power of 900) onboard the Hubble Space Telescope: “The flux at optical wavelengths corresponds to V ∼ 18.5. The spectra show many narrow, unresolved emission lines, as well as broad, blueshifted emission features. Also present are emission bands of Fe II at 225–300, 410–470, and 510–540 nm that are often seen in QSO spectra (Francis et al. 1991, Ap.J. 373, 465). Narrow lines include [O III] (500.7, 495.9, 436.3 nm), [Ne III] (386.9 nm), [Ne V] (342.5 nm), [C III (190.9 nm), [C IV (154.9 nm), and [N IV (148.8 nm). The most prominent broad line is Mg II (280 nm), with a total flux of 1.2 × 10^{13} erg cm^{−2} s^{−1}. Mg II has a triangular profile that rises sharply beginning at −6500 km/s, peaks at −3800 km/s, and then declines to zero flux at ∼ +4000 km/s. Other broad emission is seen in Lyα, N V (124 nm), O I (130 nm), Si IV (138 nm), Hβ, and an unidentified line observed at 391 nm. The broad features are all shifted by −4500 km/s and have widths of 3000–4000 km/s FWHM except for Hβ, which has a width of only 1200 km/s. These spectra confirm that SN 1998S is colliding with circumstellar gas (cf. IAUC 6830), and this interaction should be observable at other wavelengths.”

CORRIGENDA On IAUC 7046, ‘Supernovae 1998eo, 1998ep, 1998eq’, the offsets in

the table were given incorrectly (though the accompanying text is correct). The offset for SN 1998eo should read near center. The offset for SN 1998ep should read 6′′ W, 3′′ S.

COMET C/1995 O1 (HALE-BOPP)

Ephemeris extension to IAUC 7018 (elements MPC 32410):

1998/99 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Nov. 23 7^{h}01.^{m}9 −71^{o}42.^{′}0 7.057 7.000 82.^{o}6 8.^{o}0 10.7
Dec. 3 6 46.4 −72 43.6 7.141 7.084 82.7 7.9 10.8
13 6 28.0 −73 26.2 7.225 7.168 82.8 7.8 10.8
23 6 08.0 −73 47.6 7.310 7.252 82.8 7.7 10.9
Jan. 2 5 48.0 −73 47.4 7.395 7.335 82.7 7.6 11.0
1998 November 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green